7. Identify the reagent(s) that you would need to accomplish the following transformations. -O
Answer #1

(1) Reagent - H2O , H+ (lewis acid as catalyst)

this reaction is called hydration of alkenes

The mechanism of hydration involves electrophlic addition of the proton to the double bond to form a carbocation intermediate. Addition of water in the second step results in formation of an oxonium ion, which, upon deprotonation, gives the alcohol.

RA =CH2 + H2O H* = R + CH3 HOL -H H-0+ CH3 CH RR RR

when the carbocation is formed, rearrangement is possible. In the given problem, rearrangement occurs from 20 to 30 carbocation so -OH group is substituted at benzene ring.

(2) Reagent - (i) BH3 ,THF (ii) H2O2 , OH-

This reaction is called hydroboration–oxidation reaction

HC нВН, НО, BH3 нон н,с-с-сн, c=cн, нс - Н-с-с-сн, CH ? OH CH3

The hydroxyl group attaching to the less-substituted carbon. The reaction thus provides a complementry product to the above reaction. In your problem -OH gets attached to 1 degree carbon instead of 3 degree.